OpenStudy (anonymous):

Need help graphing the intersection of two lines and naming the type of system. Line y1= (1/Pix-4) + (Sqrt(2)/2.1x) Line y2 = 1.3

OpenStudy (anonymous):

can you rewrite the lines using parenthesis?

OpenStudy (anonymous):

Is the format wrong?

OpenStudy (anonymous):

\[\frac{ 1 }{ \Pi(x) - 4 } + \frac{ \sqrt{2} }{ 2.1(x) }\]

OpenStudy (anonymous):

That is what y1 looks like, if that helps.

OpenStudy (anonymous):

intersection would just be y1=y2 solve for x

OpenStudy (anonymous):

as for graphing, y2 is just a line make a chart for y1 or use a graphing calculator seeing how hte numbers are a bit ridiculous

OpenStudy (anonymous):

Line y1= (1/Pix-4) + (Sqrt(2)/2.1x) Line y2 = 1.3 The two lines intersect when they are equated together. so... 1.3 = (1/Pix-4) + (Sqrt(2)/2.1x) 2.73x = ((2.1x)/(pix-4)) + sqrt(2) 2.73x = 2.1x + (pix -4)(sqrt(2)) .63x = pix(sqrt(2) - 4sqrt(2) -3.812882938x = -4sqrt(2) x = 1.483616031 i used the whole value of pi. y is given

OpenStudy (anonymous):

or you can use the graphing calculator.

OpenStudy (anonymous):

My study plan is requesting two values of x. I tried an emulator calculator, ti-83, and it is producing a "heartbeat" intersecting at more than one point, so I think I am entering it wrong. franciscanmonk, your solution looks really accurate, but is there another x value?

OpenStudy (anonymous):

x=1.8 x=.351